OpenStudy (anonymous):

Antony is 30 years older than his daughter. In five years time, Antony will be 4 times as old as his daughter. How old is Antony now?

OpenStudy (anonymous):

I got this *like a boss* let antony = x years old daughter = y years old we can find 2 equations y+30 = x x+5 = (y+5)*4 now solve the system by plugging x's value into the 2nd equation: (y+30) +5 = 4y +20 y -4y = 20 - 35 -3y = -15 y = 5 now we know that his daughter is 5 years old since antony is 30 years older, he is 35 (x=35)

OpenStudy (anonymous):

OH Thank you so much, I see what I've done wrong now (:

OpenStudy (anonymous):

no probs :)